OpenStudy (anonymous):

What is the arc measure of a sector with area 38.485 ft2 and radius 7?

OpenStudy (anonymous):

If the radius is 7, the entire circle would have area pi*7^2 = 49*pi =~ 153.938 But we only have 38.485, so divide that by (49*pi) and we have: 38.485 / (49*pi) =~ 0.25, that's almost exactly 1/4 of a circle (90 degrees or pi/2 radians)

OpenStudy (anonymous):

thank